Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-2x + 6y = -12
x + 3y = -3

Answer by ewatrrr(24785)

 
Hi,
-2x + 6y = -12
x + 3y = -3
-2x + 6y = -12
2x + 6y = -6
12y = -18
y = -18/12
y = -3/2 and x = 3/2 (x = -3*-3/2 -3 = 3/2
|Yes! Ordered pair Pt(1.5,-1.5)is the solution for this INDEPENDENT system:See below
Lines intersect at Pt(-2,-4)

A consistent system has at least one solution.
An inconsistent system has no solutions.
An independent system has exactly one solution. ***
A dependent system has infinitely many solutions.